Section 19  History of Mathematics

Section 19: History of Mathematics

Invited Lectures

Monday, 24 August  MA 004

14:00 – 14:45 Jeremy John Gray, The Open University, Milton Keynes, England The Riemann-Roch Theorem and Geometry, 1854-1914 (I: p.208) Chair: Umberto Bottazzini

15:00 – 15:45 Karine Chemla, CNRS and Université Paris 7, Paris, France History of Mathematics in China: A factor in World History and a Source for New Questions

(I: p.206)

Chair: Sergei S. Demidov

16:15 – 17:00 Joseph Dauben, The City University of New York, New York, USA Marx, Mao and Mathematics: the Story of Nonstandard Analysis in China and the Cultural Revolution (I: p.208) Chair: Ivor Grattan-Guiness

Short Communications and Poster Sessions

Short Communications  Monday, 24 August
Non-European Mathematics  H 2013

Chair: Shirong Guo

17:15 – 17:30 Shirong Guo, The Needham Research Institute, Cambridge, Great Britain Yang Hui’s Influence on the Mathematical Tradition of the Ming Dynasty in China (p.367)

17:30 – 17:45 Osamu Takenouchi, Osaka International University, Toyanoka-shi, Japan Summation of a sequence of natural numbers (p.369)

17:45 – 18:00 Abdelkhalek Cheddadi, École Mohammadia d’Ingénieurs, Rabat, Morocco Un problème d’arithmétique et sa résolution par le mathématicien marocain al-Hassar (XIIè s.)

(p.365)

Short Communications  Tuesday, 25 August
19th Century  H 2013

Chair: Herbert A. David

14:00 – 14:15 Mircea Radu, University Bielefeld, Bielefeld, Germany Justus Grassmann’s Account of Synthesis in Mathematics: between Schelling and Kant (p.368)

14:15 – 14:30 Heiner Lichtenberg, Bundesministerium der Finanzen, Bonn, Germany The decoding of the Callippian Cycle in the Gregorian Calendar for the period between 1900 and 2099 (p.367)

14:30 – 14:45 Herbert A. David, Iowa State University, Ames, Iowa, USA Early sample measures of variability (p.366)

ICM’98 – Program  1

History of Mathematics  Section 19

Data Bases  H 2013

Chair: Patrick D.F. Ion

15:00 – 15:15 June E. Barrow-Green, The Open University, Milton Keynes, UK Mathematics in Britain 1860-1940: a database for the World Wide Web (p.365)

15:15 – 15:30 Harry B. Coonce, Mankato State University, Mankato, MN, USA The Genealogy Project for Mathematicians (p.365)

15:30 – 15:45 Patrick D.F. Ion, Mathematical Reviews, Ann Arbor, MI, USA The symbolism of mathematics: from pasigraphy in 1897 to MathML in 1998 (p.367)

Short Communications  Wednesday, 26 August
Miscellaneous  H 2013

Chair: Seemanapalli V. Sastry

16:15 – 16:30 Seemanapalli V. Sastry, Federal University of Paraiba, Paraiba, Brazil History of Mathematics: Yesterday, Today, and Tomorrow (p.368)

Poster Session  Wednesday, 26 August 14:00 – 15:45  MA 144

Carlos Gomez Bermudez, Universidade da Coruǹa, Sada, Spain

Le système des nombres chez G. Cantor (p.366) Harald Gropp, Universität Heidelberg, Heidelberg, Germany

Levi, Steinitz, Zacharias, and configurations — Three German mathematicians in the 20^{th} century and a combinatorial structure (p.366)

Aleksandar T. Lipkovski, University of Belgrade, Belgrade, Yugoslavia

Rational curves with elliptic arcs — Serret’s solution of Legendre’s problem (p.368) Man-Keung Siu, University of Hong Kong, Hong Kong, China, Peoples Republic of

On the myth of an ancient Chinese theorem about primality (p.369)

2  ICM’98 – Program